A chemist prepares a solution of zinc oxalate ZnC2O4by measuring out 1.31mg of zinc oxalate into a 150.mLvolumetric flask and filling the flask to the mark with water.Calculate the concentration in /molLof the chemist's zinc oxalate solution. Be sure your answer has the correct number of significant digits




1
Expert's answer
2021-12-09T03:23:01-0500

M(ZnC2O4)=65,39+2x12,011+4x16,994=157,388 g/mol

In 1,31 mg=0.00131 g ZnC2O4 contains 0.00131/157,388=8,3x10-6 mol.

So concentration in solution is 8,3x10-6/0.15= 5,5x10-5 mol/L